A car dealer offers you a choice of 0% financing for 60 months or $2500 cash back on a new vehicle. You have a pre-approved 60-month loan you can use from your credit union at a 4% interest rate. If the monthly payments at 0% are $16.67 per $1000 financed, and the monthly payments at 4% are $18.41 per $1000 financed, what is the range of new car prices for which the cash back option will cost you less? For what range of car prices should you take the 0% financing?
This is what I have..
$16.67 * $1000 = $16,670.
$16,670. /60 = $277.83
$18.41 * $1000 = $18,410. - $2500 = $15,910.
$15,910 /60 = $265.17

Answer by Theo(3458)

The problem states that if you take the 0% financing, then you don't get $2500 cash back, but if you take the 4% financing, then you do get $2500 cash back.

A basic assumption is that you use the $2500 cash back as a down payment on the car.

This means that if you take the 4% financing with the $2500 cash back, then the amount you have to finance is $2500 less.

Let x = the price of the car.

With 0% financing, you have to finance the full price of the car which is equal to x.

With 4% financing, you have to finance $2500 less because you are given $2500 Cash Back which you use as a down payment on the car.

this means you have to finance (x - 2500).

You have 2 equations to work with.

Your first equation with the 0% financing is:

total payments = 60 * (16.67 / 1000) * x

Your second equation with the 4% financing is:

total payments = 60 * (18.41 / 1000) * (x - 2500)

You want to know when the 0% financing plan is better than the 4% financing plan.

It is better when the total payments under the 0% financing plan are less than the total payments under the 4% financing plan.

You find this by setting up an equation that compares the total payments under the 0% financing plan with the total payments under the 4% financing plan.

You get:

Total Payments with 0% financing plan < Total Payments with 4% financing plan.

Replace the above statement with the equations for each of those options to get:

60 * (16.67 / 1000) * x < 60 * (18.41 / 1000) * (x - 2500)

You solve this equation in the following manner:

Divide both sides of this equation by 60 and multiply both sides of this equation by 1000 to get:

16.67 * x < 18.41 * (x-2500)

Simplify this equation by multiplying out the right hand side of the equation to get:

(16.67 * x) < (18.41 * x) - (18.41 * 2500)

Simplify this further to get:

(16.67 * x) < (18.41 * x) - 46025.

Add 46025 to both sides of this equation and subtract 16.67 * x from both sides of this equation to get:

46025 < (18.41 * x) - (16.67 * x)

Simplify this further to get:

46025 < 1.75 * x

Divide both sides of this equation by 1.74 to get:

26451.14943 < x

This is the same as:

x > 26451.14943

Since x is the price of the car, what this means is that you will be better off with the 0% financing option when the price of the car is greater than $26,451.14943.

In order to confirm that this value is good, we need to put in some values for the price of the car to see when the 0% financing plan is better and when the 4% financing plan is better.

If the price of the car is less than $26,451.14943, the 0% financing plan should cost more (total payments will be higher).

If the price of the car is greater than $26,451.14943, the 0% financing plan should cost less (total payments will be lower).

Your break even point should be when the price of the car is exactly $26,451.14943.

I will create a table assuming the price of the car is either:

$20,000 (less than $26,451.14943)
$26,451.14943 (EQUAL TO $26,451.14943)
$30,000 (greater than $26,451.14943)

The table is shown below:

POC = Price of car.
TP0 = Total Payments with the 0% financing plan.
TP4 = Total Payments with the 4% financing plan.

POC                     TP0                     TP4	          Result

$20,000.00		$20,004.00		$19,330.50        TP0 > TP4
$26,451.15		$26,456.44		$26,456.44        TP0 = TP4
$30,000.00		$30,006.00		$30,376.50        TP0 < TP4


From the above Table:

When the price of the car is less than $26,451.15, the 0% financing plan is more expensive.

When the price of the car is equal to $26,451.15, the 0% financing plan and the 4% financing plan cost the same. This is the break even point.

When the price of the car is greater than $26,451.15, the 0% financing plan is less expensive.

These numbers confirm the results of the equation analysis.

The formulas used are:

Total Payments with the 0% financing plan = (60 * 16.67)/1000 * x

Total Payments with the 4% financing plan = (60 * 18.41)/1000 * (x-2500)

x is the price of the car.

With the 0% financing plan, you have to finance x.

With the 4% financing plan, you have to finance (x - 2500).

This is because you got $2500 cash back with the 4% financing plan and you used that $2500 as a down payment on the car which meant that you had to finance $2500 less.